OpenStudy (anonymous):

ln7x^2 find the derivative

OpenStudy (anonymous):

y=ln7x^2=ln7+2lnx after differentiating ... hence y'=2/x

OpenStudy (anonymous):

y'=1/7x^2 times derivative of 7x^2 =14x y'= 14x/7x^2 y'= 2/x